OpenStudy (anonymous):

Solve the rational equation: 2/x+3/(x-1)=23/4x

OpenStudy (anonymous):

\[2/x+3/(x+1)=23/(4x)\] \[(4/4)(2/x) +3/(x+1)=23/(4x)\] Get similar values on both sides for combining: \[8/(4x)+3/(x+1)=23/(4x)\] Move similar values to same side: \[3/(x+1)=15/(4x)\] Cross multiply: \[12x=15(x+1)\] \[12x=15x+15\] Combine simlar values: \[-3x=15\] \[x=-5\]

OpenStudy (anonymous):

awesome! Thank you so much, I was going about it all wrong.

OpenStudy (anonymous):

only my answer says it should be +5 not -5

OpenStudy (anonymous):

oh, cause I wrote the problem wrong ok :3 well thank you!

OpenStudy (anonymous):

Np :)